1. The total production of the mobile company last week is 160,000 items. The design capacity of the factory of this company is 900 items per hour. The effective capacity of this factory is 195,000 items. The staff of this factory operate 5 days/week, 5 and 9 hours shift.
1
Expert's answer
2020-07-14T09:41:29-0400
"Q_E=(t_1+t_2)\times n \times Q_0"

"t_1= 5 h"

"t_2=9 h"


"n=5 day\/week"


"Q_0=900"


"Q_E=(5+9)\times 5 \times 900=63000"

Therefore, the design capacity provides for a different work schedule, that is, a different number of shifts and work days per week.